OpenStudy (anonymous):

The count in a bacteria culture was 700 after 10 minutes and 1500 after 35 minutes. Assuming the count grows exponentially, what was the initial size of the culture? / Find the doubling period. / find the population after 80 minutes. / When will the population reach 14000?

OpenStudy (kropot72):

The size of the culture as a function of time can be expressed as: \[\large y=Ae^{kt}\ .........(1)\] where A is the initial size and k is the growth constant. Plugging the given values into equation (1) gives: \[\large 700=Ae^{10k}\ ...........(2)\] and \[\large 1500=Ae^{35k}\ .........(3)\] Dividing (3) by (2) we get: \[\large \frac{1500}{700}=\frac{Ae^{35k}}{Ae^{10k}}\ ...........(4)\] Equation (4) simplifies to: \[\large \frac{15}{7}=e^{25k}\ .............(5)\]

OpenStudy (kropot72):

Taking natural logs of both sides of (5) we get: \[\large \ln \frac{15}{7}=25k\ .............(6)\] and solving (6) gives k = 0.0305 Substituting the value of k into (2) we find that the initial size of the culture A = 516.

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

You're welcome :)